OpenStudy (anonymous):

Simplify the expression. [90 – (12 + 13)](15 ÷ 3)

OpenStudy (anonymous):

@some.random.cool.kid

OpenStudy (anonymous):

[ 90 - (12 + 13) ] (15 / 3) = [ 90 - 25] (15/3) = 65 ( 5) = 325

OpenStudy (some.random.cool.kid):

yep correct

OpenStudy (some.random.cool.kid):

btw i was afk doing hard math

OpenStudy (anonymous):

Which could be the next step in simplifying this expression? 24 + (6 + 29) A. 24 • 6 + 24 • 29 (use the distributive property) B. 24 • 6 + 29 (use the associative property) C. (24 + 6) + 29 (use the associative property) D. 24 + 6 • 29 (use the commutative property)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Which could be the next step in simplifying this expression? 24 + (6 + 29) A. 24 • 6 + 24 • 29 (use the distributive property) B. 24 • 6 + 29 (use the associative property) C. (24 + 6) + 29 (use the associative property) D. 24 + 6 • 29 (use the commutative property)

OpenStudy (anonymous):

@some.random.cool.kid

OpenStudy (some.random.cool.kid):

back

OpenStudy (some.random.cool.kid):

most likely c

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

yep...its C

OpenStudy (anonymous):

let me get a question so just hold on a minute

OpenStudy (anonymous):

Wind is related to the movement of warm and cold air masses. Which kind of heat transfer does this represent? A. radiation B. conduction C. convection D. absorption

OpenStudy (niahloves):

C